Kafi and Rohan take part in a running competition. Kafi's speed was 15 ms^-1 after 10 seconds of start of the competition, at which time Rohan was 50 meter behind Kafi. In this case, if they both finish the competition at the same time, then their momentum is equal when they reach the finish line. What is the ratio of their mass?
1
Expert's answer
2020-06-01T14:26:22-0400

"v_R+\frac{s}{t}=v_K"

"m_Rv_R=m_Kv_K"

"\frac{m_R}{m_K}=\frac{v_K}{v_K-\frac{s}{t}}=\frac{15}{15-\frac{50}{10}}=1.5"